Last month sellers of good Y took in $100 in total revenue on sales of 50 units


1
Expert's answer
2021-11-22T11:04:44-0500

The price of good Y increased from "\frac{100}{50}=" $2 to "\frac{120}{40}=" $3. Meanwhile, the quantity demanded of good X increased from 20 to 40.

Thus, the Cross price elasticity is:

"CPED=\frac{\frac{(40-20)}{(40+20)}}{\frac{(3-2)}{(3+2)}}=\frac{5}{3}=1.67"

Since the cross price elasticity of demand is positive, the two goods are substitutes.

They are substitutes and have a cross price elasticity of 1.67.